I need some help on this please.
I know that "technically" tese do not "satisfy" the first set but it does for the second, however I feel like I am missing some steps

Is (-2,3) a solution for 3x+y =2 and x-2y=-8?
3(-2)+ 3 = 2
-6+3=-3 therefor no it is not
-2-(2)(3) = 8
(-2) - (6) = 8 Yes it is a solution for this one...
What am I missing?




Found 2 solutions by Boreal, greenestamps:
Answer by Boreal(15235)

3x+y=2
x-2y=-8
6x+2y=4
7x=-4
x=(-4/7)
for y, -12/7+ y=14/7
y=26/7
check in second
-4/7-52/7=-56/7=-8
(-2, 3) is not a solution
-6+3 is not 2

Answer by greenestamps(13200)



The two equations are of straight lines, so they will only intersect in one place.

The other tutor answered your question by solving the pair of equations to find the point of intersection of the two lines; since that point was not (-2,3), his answer was "no".

That is way too much work to answer the question. To see if a given point is a solution to both equations, you only need to plug the coordinates of the given point into both equations.

The specific confusion you seemed to express in your message was because the given point satisfied the second equation but not the first. But the answer to the question is "yes" only if it satisfies BOTH equations.

Since the given point satisfies only one of the two equations, the answer to the question is "no".